OpenStudy (anonymous):

The complete solution set of Y for |x-1| + |x-2|+|x-4| > 10 is ...

OpenStudy (anonymous):

is it (-\[is \it (-\infty,0) U (5,\infty)\]

OpenStudy (anonymous):

\[(-\infty,0) U (5,\infty)\]

OpenStudy (anonymous):

\[(-\infty,0) U (5,\infty)\]

OpenStudy (anonymous):

x=5 gives 4+3+1 = 8 not 10.

OpenStudy (anonymous):

The point is near 5 though.

OpenStudy (anonymous):

I'm looking at a computer generated graph so I can't tell you exactly what it is yet though.

OpenStudy (anonymous):

Its correct mate. Your answer is right!

OpenStudy (anonymous):

(-inf,-1) is part of the set.

OpenStudy (anonymous):

(-inf,-1) U (17/3,inf)